OpenStudy (anonymous):

Jan purchased 3 1/3 yards of yellow fabric. She also purchased 1½ yards of white fabric and ¾ yard of purple fabric. What is the total amount of fabric that Jan purchased?

OpenStudy (anonymous):

\[3\frac{1}{3} + 1\frac{1}{2} + \frac{3}{4} = \frac{10}{3} + \frac{3}{2} + \frac{3}{4} = \\ \frac{4(10)}{12} + \frac{6(3)}{12} + \frac{3(3)}{12} = \frac{40}{12} + \frac{18}{12} + \frac{9}{12}\]\[\frac{67}{12}\] or \[5\frac{7}{12}\] as a mixed number